The combustion of coal with oxygen forms CO2 and releases 94 kcal of energy.
C(s) + O2(g)

CO2(g)     
Δ
H = −94 kcal/mol
How much energy is released when 23.0 g of C reacts?
&times
10(?)
kcal of energy released
1
Expert's answer
2019-05-24T01:59:52-0400

C(s) + O2(g) → CO2(g), ΔH = −94 kcal/mol


One mole of carbon liberates 94 kcal of heat when it undergoes complete combustion.

Calculating amount of carbon:

"n(C)=\frac{m(C)}{M(C)}=\frac{23.0g}{12\frac{g}{mol}}=1.92mol"

Thus the given amount of carbon will generate

"Q=-\Delta H \cdot n(C) = 94\frac{kcal}{mol} \cdot 1.92 mol = 180.5 kcal = 1.805 \times 10^2 kcal"